OpenStudy (anonymous):

factor the trinomial completely... x^2+2x-24

OpenStudy (cwrw238):

you need to find 2 numbers a and b where ab = -24 and a - b = +2 the factors will be (x +a)(x + b)

OpenStudy (anonymous):

Either you can go the long route of using the quadratic formula or you can look at the 2x and the -24 and try and think about what two numbers would equal -24 if you multiplied them, but would also equal 2 if you added them

OpenStudy (cwrw238):

* a + b = 2

OpenStudy (cwrw238):

Clue: 6 * 4 = 24 but we need -24

OpenStudy (cwrw238):

so either the 2 or the 6 is negative we need the addition to be + 2 so which one will be negative?

OpenStudy (anonymous):

I think you mean either the 4 or 6 is negative

OpenStudy (cwrw238):

oh sorry - 4 or 6